### v1

# A Diffusion-based Xray2MRI Model: Generating Pseudo-MRI Volumes From one Single X-ray

Abstract
Knee osteoarthritis (KOA) is a prevalent musculoskeletal disorder, and X-rays are commonly used for its diagnosis due to their cost-effectiveness. Magnetic Resonance Imaging (MRI), on the other hand, offers detailed soft tissue visualization and has become a valuable supplementary diagnostic tool for KOA. Unfortunately, the high cost and limited accessibility of MRI hinder its widespread use, leaving many patients with KOA reliant solely on X-ray imaging. In this study, we introduce a novel diffusion-based Xray2MRI model capable of generating pseudo-MRI volumes from one single X-ray image. In addition to using X-rays as conditional input, our model integrates target depth, KOA probability distribution, and image intensity distribution modules to guide the synthesis process, ensuring that the generated corresponding slices accurately correspond to the anatomical structures. Experimental results demonstrate that by integrating information from X-rays with additional input data, our proposed approach is capable of generating pseudo-MRI sequences that approximate real MRI scans. Moreover, by increasing the inference times, the model achieves effective interpolation, further improving the continuity and smoothness of the generated MRI sequences, representing one promising initial attempt for cost-effective medical imaging solutions.
Keywords Knee osteoarthritis  $\cdot$ X-ray  $\cdot$ Magnetic Resonance Imaging  $\cdot$ Diffusion-based
1 Introduction
Knee OsteoArthritis (KOA) is a progressive disease marked by the degeneration and damage of articular cartilage, changes at the joint margins, and reactive hyperplasia of the subchondral bone [1]. Various factors, such as age, obesity, stress, and trauma, can contribute to its development [2]. Those affected often endure severe pain and limited mobility, which can greatly diminish their quality of life and elevate the risk of chronic conditions like cardiovascular disease [3]. Despite extensive research, the precise cause of KOA remains unknown, and there is currently no cure [4]. Traditionally, X-ray imaging has been the most common method for diagnosing KOA due to its accessibility and ability to reveal changes in bone structure, such as joint space narrowing and osteophyte formation [5]. In 1957, Kellgren and Lawrence introduced the Kellgren-Lawrence (KL) grading system for assessing KOA [6] on X-ray. According to Table 1, this system categorizes KOA into five distinct grades, determined by the presence and severity of symptoms such as osteophytes and Joint Space Narrowing (JSN).
Table 1: Detailed Overview of the KL Grading System
Grade Severity Detailed description
KL-0 None Definitive absence of any osteoarthritis signs
KL-1 Doubtful Possible presence of initial osteophytic lipping
KL-2 Minimal Certain osteophytes formation and potential JSN
KL-3 Moderate Multiple moderate osteophytes, confirmed JSN, some
bone sclerosis, and potential bone end deformities
KL-4 Severe Large and numerous osteophytes, confirmed JSN, and
definitive deformation of bone ends
However, X-ray imaging has limitations, particularly in visualizing soft tissue structures and early cartilage changes [7]. Magnetic Resonance Imaging (MRI), on the other hand, provides a more comprehensive assessment of the knee joint by offering detailed images of both bone and soft tissues, including cartilage, menisci, and synovium [8]. MRI can detect early signs of KOA that are not visible on X-rays, such as bone marrow lesions, synovitis, and early cartilage degeneration [9]. However, the widespread adoption of MRI in the clinical management of KOA is hindered by several challenges, most notably the high cost associated with MRI scans [10]. The expense of MRI technology, including the costs of the equipment, maintenance, and specialized personnel, makes it less accessible compared to traditional X-ray imaging [11]. This economic barrier is particularly significant in resource-limited settings and for patients without adequate health insurance coverage. In addition to the high financial cost, MRI scans require more time and patient cooperation, as they necessitate the patient to remain still for extended periods [12], which can be difficult for those experiencing severe pain or limited mobility due to KOA.
To address these limitations, researchers have explored various techniques to enhance MRI reconstruction and reduce its dependency on traditional methods. Fessler et al. [13] provides an in-depth review of iterative algorithms used for improving MRI image reconstruction, particularly in scenarios where traditional inverse Fourier transforms are inadequate. The author highlights the limitations of conventional methods, especially when dealing with non-Cartesian sampling, under-sampled data, or the presence of nonlinear magnetic fields, and then focuses on model-based approaches, which utilize iterative methods that incorporate physical models of MRI acquisition, allowing for better image quality by addressing inhomogeneity and under-sampling issues. The authors demonstrate that model-based methods can significantly enhance image reconstruction, providing superior results compared to traditional techniques, though often at the cost of increased computation. In [14], Han et al. proposed leveraging the k-space to improve the efficiency and accuracy of MRI reconstruction. The authors introduced a framework inspired by the low-rank Hankel matrix completion approach, which directly interpolates missing k-space data and enhances the representation of MRI data by learning the underlying signal structure, resulting in high-quality images even from undersampled data. The proposed k-space deep learning approach significantly reduces the computational complexity and memory requirements. In [15], Lustig et al. leveraged compressed sensing to significantly accelerate MRI scan times without compromising image quality. The authors explore the inherent sparsity of MR images, both in the image domain and in transform domains such as spatial finite differences or wavelet coefficients. By randomly undersampling k-space and using nonlinear recovery schemes, the study shows that high-quality MR images can be reconstructed from incomplete data. This approach effectively reduces acquisition time and mitigates artefacts that arise from undersampling. The authors conclude that compressed sensing has the potential to substantially reduce scan times in clinical settings, making MRI more efficient and accessible.
With the rapid advancements in deep learning technologies, the field of MRI reconstruction has seen significant growth [16] [17] [18]. Recently, combining multi-modal and diffusion-based models has become a new trend for image generation. In [19], Zhan et al. presented a novel medical generative model, MedM2G, designed to unify multiple medical modalities (e.g., CT, MRI, X-ray) and tasks (e.g., text-to-image, image-to-text, modality translation) into a single, efficient framework. The authors address the challenge of aligning and generating medical multi-modal data, particularly in the absence of large, well-paired datasets. They propose a central alignment strategy to efficiently align different modalities using text as the central modality and introduce medical visual invariant preservation to retain key clinical knowledge from each modality. The model leverages a multi-flow cross-guided diffusion process that facilitates adaptive interaction among modalities, enhancing the generation of accurate and high-quality medical images. In [20], Jiang et al. introduced a novel diffusion-based approach to multi-modal MRI synthesis. Traditional diffusion models for MRI synthesis typically operate in the original image domain, leading to high memory consumption and less effective handling of anatomical structures. CoLa-Diff addresses these challenges by operating in the latent space to reduce memory overhead while preserving anatomical details. The model introduces structural guidance using brain region masks and an auto-weight adaptation mechanism to balance multiple input modalities effectively during synthesis. The authors demonstrate that CoLa-Diff outperforms existing state-of-the-art models in multi-modal MRI synthesis tasks, showing superior results in terms of PSNR and SSIM. In [21], Kim et al. propose an advanced method for translating MRI images between different modalities. The authors introduce an Adaptive Latent Diffusion Model (ALDM) that leverages a novel multiple-switchable spatially adaptive normalization (MS-SPADE) block to transform source latent representations into target-like latent. This approach enables high-quality image synthesis across multiple modalities without requiring patch cropping, preserving the global information of the images. By conditioning the model with target modalities and applying the ALDM in a compressed latent space, the proposed method significantly reduces computational complexity while achieving superior results in both one-to-one and one-to-many image translation tasks. The authors highlight the model’s effectiveness in generating anatomically accurate and high-fidelity 3D medical images, outperforming existing methods in both qualitative and quantitative evaluations.
Inspired by the above work, we propose a novel approach named Xray2MRI, based on the Conditional Latent Diffusion Model (CLDM) [22]. Specifically, during the training phase, each real MRI slice is first noised and then progressively denoised using the corresponding X-ray as conditional input and the adaptive-weight guidance of the target depth, image intensity distribution, and KOA probability distribution to reconstruct the input MRI slice. In the inference phase, given a knee X-ray and a target depth, the model progressively denoises from Gaussian noise to generate the corresponding MRI slice at that depth. Then, the target depth is adjusted and the denoising process is repeated iteratively for each depth until all the required slices are synthesized. These individual slices are then stacked together to form a full 3D MRI volume. It is noteworthy that the reconstructed MRI is not intended to serve as a clinical replacement for a real MRI. Still, it marks a pioneering cross-modal attempt to generate corresponding pseudo-MRI volumes from X-rays, for seeking cost-effective medical imaging solutions.
The primary contributions of this paper include the following:
  • $\bullet$
    A novel CLDM-based framework, Xray2MRI, was proposed to generate the pseudo-MRI volumes from a single X-ray image.
  • $\bullet$
    In addition to the target depth, information about the KOA probability distribution and image intensity distribution from the X-ray is combined via an adaptive-weight module to guide the denoising process.
  • $\bullet$
    A new SSIM-based evaluation metric was specially developed. The multi-indicator assessment approach ensures both quantitative and qualitative validation of the synthesized MRI images for both image quality and clinical relevance.
  • $\bullet$
    The model showcases effective interpolation between slices by increasing the inference times (i.e., exceeding the depth of the original MRI volume), resulting in smoother transitions and improved image continuity.
2 Proposed approach
2.1 Classical conditional latent diffusion model
Before delving into the proposed approach, we provide a brief overview of the classic conditional latent diffusion model, which serves as the foundation for our methodology.
Figure 1: The structure of the classical conditional latent diffusion model.
Figure 2: The flowchart of the proposed approach, with black arrows indicating data flow, starts by extracting a slice at position $d$ from a real MRI sequence, denoted as $x_{d}$. This slice is encoded by the encoder $\mathcal{E}_{1}$ to produce a latent representation $z_{d}$, which is progressively noised following the distribution $q(z_{t}|z_{t-1})$, ultimately generating the noised latent representation $z_{T}$. The corresponding X-ray image $x_{c}$ is then used as a conditional input, encoded by $\mathcal{E}_{2}$, and concatenated with $z_{T}$ to form the initial input for the downsampling network. During the denoising process, the target depth $d$ is embedded at various stages of the downsampling phase. Concurrently, the KOA probability distribution $p$, obtained from a pre-trained KOA classifier $\mathcal{C}$, and the image intensity distribution $i$ from the X-ray image are integrated into the upsampling phase through an adaptive weighting module, providing dynamic guidance. This denoising process is repeated for $T-1$ iterations to progressively refine the latent representation, resulting in the denoised latent $\hat{z}_{d}$, which is decoded by $\mathcal{D}$ to reconstruct the MRI slice $\hat{x}_{d}$ at depth $d$. During inference, the process begins with $z_{T}$ and is repeated $c$ times, corresponding to the number of channels in the original MRI sequence. Finally, the full pseudo-MRI sequence $\hat{x}$ is obtained by stacking all the generated MRI slices in sequential order (i.e., $\{\hat{x}_{d=1}$, $\hat{x}_{d=2}$,…,$\hat{x}_{d=c}\}$).
As shown in Fig. 1, a conditional latent diffusion model is a type of generative model used for producing data (such as images, audio, or text) based on certain conditions or inputs. This model combines the strengths of diffusion models and latent variable models, allowing it to generate complex data distributions while being guided by specific conditions. Different modules are to be introduced, respectively.
2.1.1 Encode process
The latent variable model is used to map the input image $x$ to a latent space using an encoder:
$z=\mathcal{E}(x)$ (1)
where $\mathcal{E}$ is an encoder which maps an image $x$ to a latent representation $z$.
2.1.2 Diffusion process
A diffusion process describes the evolution of a random variable over time. Through the forward diffusion process, for the original latent representation $z_{0}\sim q(z_{0})$, each step of the diffusion process, which comprises a total of $T$ steps, involves adding Gaussian noise $n$ to the data obtained from the previous step $z_{t-1}$ as follows:
$q(z_{t}\mid z_{t-1})=\mathcal{N}(z_{t};\sqrt[]{1-\beta_{t}}z_{t-1},\beta_{t}\mathbf{I})$ (2)
where $\beta_{t}$ represents the variance used for each step in the range of $[0,1]$, and $\mathbf{I}$ represents the identity matrix. The entire diffusion process constitutes a Markov chain [23]:
$q(z_{1}\mid z_{0})=\prod_{t=1}^{T}q(z_{t}\mid z_{t-1})$ (3)
Here, we directly sample $z_{t}$ for any step $t$ based on the original data $z_{0}$: $z_{t}\sim q(z_{t}\mid z_{0})$ defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}t=\prod{i=1}^{t}\alpha_{i}$. Through the reparameterization technique, we have:
$z_{t}$ $=\sqrt{\alpha_{t}}z_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1}$ (4)
$=\sqrt{\alpha_{t}}(\sqrt{\alpha_{t-1}}z_{t-2}+\sqrt{1-\alpha_{t-1}}\epsilon_{t-2})+\sqrt{1-\alpha_{t}}\epsilon_{t-1}$
$=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\quad\forall\epsilon_{t}\sim\mathcal{N}(0,\mathbf{I})$
2.1.3 Denoising and Sampling
After the diffusion process, denoising is performed by sampling from the updated noise distribution. Estimating distribution $q(x_{t-1}\mid x_{t})$ requires the utilization of the entire training set. Typically, a neural network, such as a U-Net [24], is employed to estimate these distributions. Here, the reverse process is also defined as a Markov chain composed of a sequence of Gaussian distributions parameterized by neural network parameters:
$p_{\theta}(z_{0:T})=p(z_{T})\prod_{t=1}^{T}p_{\theta}(z_{t-1}\mid z_{t}),\quad p(z_{T})=\mathcal{N}(z_{T};0,\mathbf{I})$ (5)
where
$p_{\theta}(z_{t-1}\mid z_{t})=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t,y),\Sigma_{\theta}(z_{t},t,y))$ (6)
$p_{\theta}(z_{t-1}\mid z_{t})$ represents parameterized Gaussian distributions, the mean $\mu_{\theta}(z_{t},t,y)$ and variance $\Sigma_{\theta}(z_{t},t,y)$ with the condition $y$ are determined by trained learning networks.
2.1.4 Training process
The encoder-decoder and diffusion modules are trained separately. For the encoder-decoder, the reconstruction loss $\mathcal{L}_{rec}$ is as:
$\mathcal{L}_{rec}=||(x,\mathcal{D}(\mathcal{E}(x))||_{2}^{2}$ (7)
The parameters $\theta$ of the diffusion module aim to minimize the distance between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}$. The diffusion loss $\mathcal{L}_{diff}$ quantifies this difference as the expected value of the squared norm between the actual noise and the noise estimated by the neural network:
$\mathcal{L}_{diff}=\mathbb{E}_{z\sim E(x),t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}(z_{t},t,c(x,y))||_{2}^{2}$ (8)
where $\mathbb{E}$ represents the expectation over the joint distribution of the original latent representation $z_{0}$ and Gaussian noise $\epsilon$.
2.2 Proposed approach
As presented in Fig. 2, our approach consists of three main modules: an auto-encoder module, a CLDM module, and a guidance module. $x_{d}$ is one slice of the original knee MRI sequence $x$ at the depth position $d$ with the size of $H\times W\times C$. $\mathcal{E}_{1}$ and $\mathcal{D}$ are encoder and decoder of the AutoencoderKL [25]. Unlike traditional autoencoders that map input data to a deterministic latent space, the AutoencoderKL introduces a probabilistic approach by encoding the input into a Gaussian distribution over the latent space. During training, the Kullback-Leibler (KL) divergence is used as a regularization term to ensure that the learned latent space closely follows a standard normal distribution, promoting smoother and more continuous transitions between latent representations. Specifically, the encoder operates at a resolution of 256$\times$256$\times$3 and begins with a convolutional layer configured with 128 channels and then proceeds through a series of stages, each with 2 residual blocks. The channel dimensions are multiplied according to the sequence 1, 2, 4, and 4. The decoder mirrors the structure of the encoder. The architecture is designed with no dropout, focusing on retaining all features without stochastic regularization during training. At the same time. the noised latent code $z_{T}$ is calculated following the Eq. 4:
$z_{T}=\sqrt{\bar{\alpha_{t}}}\mathcal{E}_{1}({x_{0}})+\sqrt{1-\bar{\alpha_{T}}}\epsilon$ (9)
Conditional image $x_{c}$ is a corresponding X-ray image, which is then encoded by the encoder $\mathcal{E}_{2}$ and concatenated with the $z_{T}$ as the input of the U-Net, serves as the beginning of the reverse process. The employed 2D U-Net architecture adopts a multi-scale hierarchical structure comprising a time embedding module, an encoder module, a middle module, and a decoder module. This UNet model takes 8-channel input and produces 4-channel output, built upon a base channel configuration of 320. The encoder module begins with a 2D convolutional layer featuring a kernel size of 3 and padding of 1, followed by a series of residual blocks with progressively increasing feature dimensions. Each residual block includes multiple convolutional layers, GroupNorm [26] for normalization, the Swish activation function [27] for non-linearity, and optional dropout for regularization. Two residual blocks are employed per downsampling stage, with attention mechanisms activated at resolutions of 4, 2, and 1. The channel count increases across network layers following a multiplier sequence of 1, 2, 4, and 4. The decoder mirrors this structure, utilizing the same residual blocks to progressively upsample feature maps, reducing the number of channels in reverse order while merging them with corresponding encoder feature maps through skip connections. Additionally, the model incorporates 8 attention heads and employs a spatial transformer with a depth of 1, conditioned on a context dimension of 768.
Next, the target depth $d$ was embedded into the downsampling stage to allow the model to condition the latent representation with spatial hierarchy early in the process. This ensures that depth-related features, such as different tissue layers and bone structures, are incorporated from the outset, preserving the anatomical coherence of the volume. In the upsampling stage, the KOA probability distribution $p$ and image intensity distribution $i$ were embedded to guide the synthesis of fine clinical features based on the spatial hierarchy established during downsampling, such as the presence of osteophytes and the size of the joint space. However, since the variation between slices in MRI is not uniform, different slices have varying levels of sensitivity to the guided information. Some slices may require more emphasis on specific anatomical features, while others may need more focus on KOA severity. To address this, we introduced an automatic weight adjustment mechanism that dynamically combines the information from the KOA probability distribution and the image intensity distribution, which ensures that the final MRI outputs maintain a higher level of accuracy and clinical relevance. The adaptive-weight module $w$ begins by independently adjusting the weights of the KOA probability distribution $p$ and the X-ray image intensity distribution $i$ using L2 norm, scaled by learnable parameters $\mu_{d}$ and $\mu_{p}$, respectively:
$G_{p}=\mu_{p}\|p\|_{2}=\mu_{p}\left(\sum_{m=1}^{h}\sum_{n=1}^{w}(p_{m,n})^{2}+\kappa\right)^{\frac{1}{2}}$ (10)
$G_{i}=\mu_{i}\|i\|_{2}=\mu_{i}\left(\sum_{m=1}^{h}\sum_{n=1}^{w}(i_{m,n})^{2}+\kappa\right)^{\frac{1}{2}}$ (11)
where $h$ and $w$ represent the height and width of the feature maps $p$ and $i$. $p_{m,n}$ and $i_{m,n}$ represent the values of the respective feature maps at the pixel located at row $m$ and column $n$. $\kappa$ is a small constant added to prevent division by zero, typically set to an empirical value of 1e-05.
Then, the calculated weights $G_{p}$ and $G_{i}$ are normalized as follows:
$\hat{G}_{p}=\frac{G_{p}}{\sqrt{G_{p}^{2}+\epsilon}},\quad\hat{G}_{i}=\frac{G_{i}}{\sqrt{G_{i}^{2}+\epsilon}}$ (12)
The KOA probability distribution $p$ and image intensity distribution $i$ are then adjusted by applying the normalized weights:
$\tilde{p}=p\left[1+\sigma(\nu_{p}\hat{G}_{p}+o_{p})\right],\quad\tilde{i}=i\left[1+\sigma(\nu_{i}\hat{G}_{i}+o_{i})\right]$ (13)
where $\sigma$ is a Sigmoid function. $v_{p}$, $v_{i}$, and $o_{p}$, $o_{i}$ are learnable scaling and offset parameters to $p$ and $i$, respectively.
To capture the interaction between $p$ and $i$, a joint condition weight $\lambda_{pi}$ is learned using an attention mechanism [28], which is represented as:
$\lambda_{pi}=\text{Attention}(p,i)=\text{softmax}\left(\frac{(W_{p}p)^{\top}(W_{i}i)}{\sqrt{p_{k}}}\right)$ (14)
where $W_{p}$ and $W_{i}$ are learnable weight matrix associated with $p$ and $i$, respectively. $\sqrt{p_{k}}$ is a scaling factor.
The final combined condition $\tilde{y}_{pi}$ with the learned joint condition weight $\lambda_{pi}$ is as follows :
$\tilde{y}_{pi}=\lambda_{pi}\cdot(\tilde{p}+\tilde{i})$ (15)
As presented before, for the diffusion process, U-Net $\mathcal{U}$ is trained based on the Eq. 8:
$\mathop{\min}_{\theta}\mathbb{E}_{z\sim\varepsilon_{1}(x_{d}),t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}\{z_{t},t,[z_{T},\mathcal{C}(x_{c}),w(p,i)]\}||_{2}^{2}$ (16)
During the reverse process, the latent variable $z_{t-1}$ is sampled using the Denoising Diffusion Implicit Model (DDIM) [29], which enables deterministic or stochastic generation depending on the chosen schedule. DDIM refines the sampling process by approximating the reverse diffusion trajectory more efficiently than traditional diffusion models. Specifically, it leverages a non-Markovian process that skips some diffusion steps, allowing for faster inference without compromising the quality of the generated images. The sampling of $z_{t-1}$ is computed as follows:
$z_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{z_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon}_{\theta}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\mathbf{\epsilon}_{\theta}$ (17)
Finally, for each reconstructed slice at the depth $d$, denoted as $\hat{x}_{d}$, it is generated through a decoder $\mathcal{D}$. During the inference phase, this process is repeated $t$ times, corresponding to the channel (depth) $c$ of the original MRI volume. The final generated pseudo-MRI sequence $\hat{x}$ is formed by stacking all generated slices, represented as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{t=c}\}$. For more clarity, Algorithm 1 describes the entire process of the proposed approach.
Input: $x$, $c$, $t$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$ and all initial parameters
Output: $w$, $\hat{x}_{d}$, $\mathcal{E}_{1}$, $\mathcal{D}$, $\mathcal{U}$ and learned parameters
while not converged do
       for $d$ in $\{0,1,...,c-1\}$ do
             Get $x_{d}\leftarrow x[d]$
             Get $\hat{x}_{d}\leftarrow\mathcal{D}(\mathcal{E}_{1}({x_{d}}))$
             Compute $\mathcal{L}_{rec}(x_{d},\hat{x}_{d})$triangleright Eq. 7
             Update Parameters of $\mathcal{E}_{1}$ and $\mathcal{D}$
       end for
end while
Freeze Parameters of $\mathcal{E}_{1}$ and $\mathcal{D}$
while not converged do
       for $d$ in $\{0,1,...,c-1\}$ do
             Get $p,i\leftarrow\mathcal{C}(x),intensity(x)$
             Get $\epsilon,z_{T}\leftarrow noise(\mathcal{E}_{1}({x_{d}}),t)$triangleright Eq. 4
             Get $\epsilon_{\theta},\hat{z}_{d}\leftarrow\mathcal{U}(z_{T},\mathcal{E}_{2}(x_{c}),t,w(p,i))$triangleright Eq. 16
             Compute $\mathcal{L}_{diff}(\epsilon,\epsilon_{\theta})$triangleright Eq. 8
             Get $\hat{x}_{d}\leftarrow\mathcal{D}(\hat{z}_{d})$
             Compute $\mathcal{L}_{rec}(x_{d},\hat{x}_{d})$triangleright Eq. 7
             Update Parameters of $\mathcal{U}$
       end for
end while
Algorithm 1 Training process of the approach
Input: $c$, $t$, $w$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$, $\mathcal{D}$ and $\mathcal{U}$
Output: $\hat{x}_{d}$ and $\hat{x}$
for $d$ in $\{0,1,...,c-1\}$ do
       Get $\epsilon\leftarrow\mathcal{N}(0,\mathbf{I})$
       Get $p,i\leftarrow\mathcal{C}(x),intensity(x)$
       Get $\hat{z}_{d}\leftarrow\mathcal{U}(\epsilon,\mathcal{E}_{2}(x_{c}),t,w(p,i))$triangleright Eq. 16
       Get $\hat{x}_{d}\leftarrow\mathcal{D}(\hat{z}_{d})$
end for
Get $\hat{x}\leftarrow\{\hat{x}_{0},\hat{x}_{1},...,\hat{x}_{c-1}\}$
Algorithm 2 Inference process of the approach
3 Experimental settings
This section provides an overview of the experimental data, the preprocessing steps undertaken, and the specifics of the experimental procedures.
3.1 Employed database
The OsteoArthritis Initiative (OAI) [30] represents a significant and rich source of data for researchers investigating KOA and related conditions. By analyzing data from 4,796 participants aged between 45 and 79 years over 96 months, the initiative provides a comprehensive longitudinal dataset. Each participant underwent nine follow-up examinations, allowing for detailed tracking of the disease’s progression and the identification of potential risk factors associated with KOA development or progression.
3.2 Data preprocessing
3.2.1 X-ray
As shown in Fig. 3, the knee joint (Fig. 3(b)) was identified from the plain radiograph (Fig. 3(a)) through the YOLOv2 learning model [31], with images sized of 299 $\times$ 299 $\times$ 1, which were resized to the size of 256$\times$256$\times$1, and image intensity distribution was normalized to [-1, 1].
(a) A standard knee plain radiograph
(b) An identified knee joint
Figure 3: A standard knee plain radiograph from the database and an identified knee joint highlighted in a red box 3(a). An identified knee joint 3(b).
3.2.2 MRI
[…]
Figure 4: Slice count distribution on each type of MRI sequences 4. Grey value distribution on each proportional position of slices within the T1-weighted MRI sequences 4.
After the processing, the slice count was fixed at 50 for the T1-weighted MRI scans as the final experimental dataset. The resulting 4,262 X-ray-MRI pairs were then randomly divided into training and validation sets with a ratio of 7:3.
3.3 Experimental details
The diffusion module utilized pre-trained weights from Zero-123 [32]. The pre-trained KOA classifier $\mathcal{C}$ and encoder $\mathcal{E}_{2}$ are sourced from the [33] and [34], respectively. During training, a base learning rate of 1e-06 was applied, with a lambda linear scheduler incorporating a warm-up period every 100 steps. The batch size was set to 64, with $T$ configured to 1,000 steps, and an Exponential Moving Average (EMA) [35] of 0.99 was used for the diffusion sampling process. The scale factor of the latent space was set as 0.2. The AdamW optimizer [36] was utilized for training over 5,000 epochs. We implemented our approach using PyTorch v1.8.1 [37] on Nvidia A100 graphics cards with 80 GB of memory.
4 Results and discussion
4.1 Comparisons with State-of-the-Art Methods
In this section, the comparison is divided into two main parts: Visualization analysis and Quantitative indicator analysis.
4.1.1 Visualization analysis
Table 2 provides a comprehensive visual comparison of corresponding pseudo-MRI slices generated by various methods, using X-ray images as input. To facilitate a clear and focused analysis, we selected three representative stages of KOA: KL-0, indicating healthy knees; KL-2, reflecting early-stage KOA; and KL-4, corresponding to late-stage KOA. The compared methods include classic LDM [38], Cola-diff [20], X-diff [39], and our proposed approach. Each row displays the X-ray input on the far left, followed by the real MRI slice and the pseudo-MRI slices generated by the respective methods. To provide a more intuitive illustration of each method’s performance, difference maps generated using a Sobel operator filter are shown beneath each slice, highlighting pixel-wise deviations from the ground truth MRI by detecting edge differences. The visual results consistently demonstrate that our model produces pseudo-MRI slices that more closely resemble the ground truth, with significantly fewer and less pronounced deviations compared to other methods. Specifically, LDM suffers from the lack of constraints provided by the input X-ray. Without this conditional guidance, the model struggles to capture the detailed structural information of the knee joint, leading to great deviations from the ground truth MRI. On the other hand, compared with LDM, although Cola-Diff and X-Diff generate bone structures that are closer to the real MRI in terms of overall shape, they both lack a significant amount of fine details. This limitation prevents them from accurately capturing and reproducing the KOA symptoms. Moreover, it is noteworthy that it is particularly evident as the KL grade of the input X-ray increases. The progressive narrowing of the knee joint space and the formation of osteophytes are learned and generated more accurately by our model, which demonstrates that our method not only ensures the generation of high-quality MRI images but also maintains clinical relevance. However, the model’s predictions in areas outside the skeletal regions are less accurate due to the absence of critical information in X-ray images such as muscle and fat distribution, which will be discussed in in Section 4.3.3.
Table 2: Visualization of the generated pseudo-MRI slices
X-ray input MRI Ground Truth LDM Cola-diff X-diff Ours
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
KL-0
Difference
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
KL-2
Difference
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
KL-4
Difference
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    To make it more intuitive and convenient, we only display here a single slice from each MRI sequence that best represents the complete knee joint.
4.1.2 Quantitative indicator analysis
Peak Signal-to-Noise Ratio (PSNR) is a measure used to assess the quality of reconstructed or compressed images compared to their original versions. The formula for PSNR is:
$\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}_{I}^{2}}{\text{MSE}}\right)$ (18)
where $\text{MAX}_{I}$ is the maximum possible pixel value of the image. MSE stands for the Mean Squared Error, which calculates the average of the squared differences between the real and synthesized images:
$\text{MSE}=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}(I(i,j)-K(i,j))^{2}$ (19)
where $I(i,j)$ and $K(i,j)$ are the pixel values at position $(i,j)$ in the real and synthesized images, respectively. $m$ and $n$ are the width and height of the image.
On the other hand, the Structural Similarity Index (SSIM) is a widely used metric for evaluating the visual similarity between two images considering human visual perception by measuring changes in luminance, contrast, and structural information. However, it does not account for the specific regions of interest (ROI) that are critical in medical imaging. In knee MRI volumes, the most relevant diagnostic information often lies within cartilage, joints, and other areas affected by disease. Therefore, in addition to SSIM, we also adopt a specially designed indicator, the so-called Region-specific Structural Similarity Index (RSSIM), to represent the clinical evaluation of bone morphology. To do so, we utilize the Canny operator with a sigma of 3 to filter and extract bone edges, followed by the application of SSIM. Specifically, we focus on the lower third of the femoral edge and the upper third of the tibial edge to capture ROIs most likely to contain areas of potential pathology, such as joint space narrowing (JSN) and osteophytes. The typical SSIM is computed as:
$\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$ (20)
where
$\begin{split}&\mu_{x}=\frac{1}{N}\sum_{i=1}^{N}x_{i},\quad\mu_{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}\\ &\sigma_{x}^{2}=\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\mu_{x})^{2},\quad\sigma_{y}^{2}=\frac{1}{N-1}\sum_{i=1}^{N}(y_{i}-\mu_{y})^{2}\\ &\sigma_{xy}=\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\mu_{x})(y_{i}-\mu_{y})\end{split}$ (21)
$\mu_{x}$, $\mu_{y}$, and $\sigma_{x}^{2}$, $\sigma_{y}^{2}$ are mean intensities and variances of the images $x$ and $y$, respectively. $\sigma_{xy}$ represents the covariance of $x$ and $y$. $C_{1}$ and $C_{2}$ serve as the stabilization constants to avoid division by zero.
[…]
Figure 5: The box plots visualize the different performance of the evaluated approaches using PSNR 5 and SSIM/RSSIM 5.
Table 3: Visualization of the synthesized intermediate frames
Starting image (real) $\stackrel{{\scriptstyle\text{Synthesized intermediate frames}}}{{\makebox[284.52756pt]{\rightarrowfill}}}$ Ending image (real)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
4.2 Impact of the inference times
For the inference times $t$, in addition to adhering to the original depth setting ($d=t=50$), we also experimented with varying inference times. The goal was to explore the impact of increasing inference times on the quality and accuracy of the generated slices. For more clarity and conciseness, we here only focus on discussing the case where $t=500$. Table 3 illustrates the synthesized intermediate frames between the original adjacent images (i.e., starting and ending images). Three distinct representative stages were evaluated: first, the transition from predominantly muscle tissue to the initial appearance of bone; second, the shift from cartilage to the developing coronal knee joint; and finally, the full visualization of the complete coronal knee joint. In each of these stages, the synthesized intermediate frames display relatively smooth and continuous transitions between bone and soft tissue structures. Importantly, the model successfully maintains key anatomical details throughout the transition, including the integrity of the bones, cartilage, and joint spaces, which underscores the robustness and effectiveness of these interpolation frames.
Moreover, the Fourier frequency domain analysis and histogram correlation measurements were conducted. As shown in Fig. 6, the intermediate frames are almost consistent with the preceding and following frames in the low-frequency region, while they contain less high-frequency information, making the overall image appear smoother, which suggests that the intermediate frames are generated through interpolation. The average histogram correlations between the original adjacent images and their four intermediate frames are 0.5404, 0.9713, 0.9334, 0.8723, and 0.5248, in sequential order. The high correlations between the consecutive intermediate frames suggest that the intensity distribution remains nearly constant throughout these frames, indicating very little variation in the pixel intensities, which points to the fact that the intermediate frames are smooth transitional frames. In contrast, the relatively lower correlations between the real start and end frames and the closest intermediate frames highlight the greater difference in intensity distributions, which suggests that the interpolation algorithm focuses on generating a gradual and seamless transition between the start and end frames, prioritizing continuity over the introduction of new or abrupt changes.
(a) Starting image
(b) Frame #$\#1$
(c) Frame #$\#2$
(d) Frame #$\#3$
(e) Frame #$\#4$
(f) Ending image
Figure 6: The Fourier frequency domain distribution for the original adjacent images and the synthesized intermediate frames between them.
The above results demonstrate that increasing the inference times $t$ significantly enhances interpolation, showcasing its potential for improving temporal resolution. By generating additional intermediate frames from a smaller set of real frames, interpolation effectively offers more continuous data without lengthening the actual scan duration.
4.3 Discussion
In this paper, we introduce a novel diffusion-based Xray2MRI model that generates full pseudo-MRI volumes from a single X-ray image, representing one promising initial step toward cost-effective medical imaging solutions. The model leverages conditional inputs from X-rays, combined with target depth, KOA probability distribution, and image intensity distribution modules, to guide the synthesis process and ensure anatomically accurate MRI slice generation. Experimental results show that the generated pseudo-MRI sequences approximate real MRI scans, delivering high image quality while preserving critical clinical relevance. Additionally, by increasing the inference times, the model exhibits effective interpolation, resulting in smoother transitions between slices. Several points are to be discussed.
4.3.1 How it works?
In this approach, we move away from focusing on the distinct imaging principles and techniques of X-ray and MRI. Instead, we only concentrate on the mapping relationship between these two imaging modalities. Much like an extremely experienced senior radiologist who, when viewing an X-ray image, could to some extent mentally infer what the corresponding MRI scans might reveal, which is our proposed AI model’s aim. To this end, for the model’s training, vast amounts of high-quality paired X-ray and MRI, along with various data augmentation techniques, have significantly propelled this.
4.3.2 Why use a 2D diffusion model instead of a 3D one directly?
We opted to use a 2D diffusion model, generating individual MRI slices at different target depths during the inference phase, which are then stacked to form a pseudo-3D MRI sequence. By using this method, the model can effectively learn and capture the spatial relationships between the individual slices, allowing it to understand inter-slice dependencies. Additionally, employing a 2D model reduces computational complexity, making both the training and inference processes more efficient compared to handling full 3D data. This approach also enhances the interpretability of the entire process, as it enables a more granular analysis of how each slice contributes to the final 3D reconstruction.
4.3.3 Strengths and limitations
This study has several notable strengths. Firstly, it introduces an innovative method for generating pseudo-MRI images from one single X-ray image, attempting to bridge the gap between these two imaging modalities using advanced AI technology. This approach represents a novel attempt in cases where MRI may not be readily available. Second, the study employs a 2D diffusion model, which not only reduces the computational resources required but also allows for more detailed control over the generated images, leading to a more interpretable and transparent reconstruction process. Finally, the study’s focus on clinically relevant metrics, including both traditional image quality assessments and pathology-specific evaluations, ensures that the generated pseudo-MRI images are not only visually similar to ground truth MRIs but also clinically relevant. This study has several limitations that warrant consideration. First, our model are highly dependent on the availability of large, high-quality paired datasets. The scarcity of such datasets, particularly in certain clinical scenarios or for rare diseases, may limit the generalizability of our approach. Second, although our model is able to approximate MRI features from X-ray inputs, it remains strongly uncertain whether these pseudo-MRI images can be reliably used for clinical diagnosis without the support of additional imaging modalities as the model cannot learn information that is not present in the input X-ray image, which is precisely the value that MRI imaging provides. Such limitation is clearly illustrated in Fig. 7. Finally, the computational resources required for training, despite being reduced by using a 2D model, high-performance hardware is still necessary, which presents a challenge for broader implementation in clinical settings.
(a) X-ray
(b) Real MRI
(c) Generated MRI
Figure 7: The X-ray 7(a) does not provide information about cartilage abnormalities, so the generated MRI 7(b) similarly lacks this detail, whereas such information is presented in the actual MRI 7(c).
4.3.4 Future work
More comprehensive information (e.g., BMI, age, gender, etc) could be effectively integrated with X-ray images to further enhance the generation performance. Moreover, to improve the model’s deployability, the use of knowledge distillation could be of interest.
5 Statements
This manuscript was prepared using data from the OAI. The views expressed in it are those of the authors and do not necessarily reflect the opinions of the OAI investigators, the National Institutes of Health (NIH), or the private funding partners.
6 Acknowledgements
The authors gratefully acknowledge the support of the Ralph Schlaeger Research Fellowship under award number 2024A018081 from Massachusetts General Hospital (MGH), Harvard Medical School (HMS).